OpenStudy (amtran_bus):

4. Determine the angle between the vectors, given that |u| |v| cos(theta) =u*v

OpenStudy (amtran_bus):



OpenStudy (amtran_bus):

the (6,3) is for the 2nd vector

OpenStudy (experimentx):

\[ \theta = \arccos \left( u \cdot v \over ||u|| \; ||v||\right)\]

OpenStudy (amtran_bus):

@Spacelimbus I done another question to de-clutter.

OpenStudy (experimentx):

\[ u = (x_2-x_1, y_2-y_1)\]

OpenStudy (amtran_bus):

vector 1 is 5.39 vector 2 is 4.12

OpenStudy (amtran_bus):

using pythag. theorem.

OpenStudy (anonymous):

plug everything back into the equation that @experimentX posted above and you're done, you already calculated the numerator in the previous exercise, now you have the length/magnitude of the vectors, thats all you need

OpenStudy (amtran_bus):

I was thinking I was going to need to do a trig function to get the angle. ok.

OpenStudy (amtran_bus):

well, duh, cos.

OpenStudy (amtran_bus):

well, invcos.

OpenStudy (amtran_bus):

is the answer ~0.89?

OpenStudy (amtran_bus):

invcos(14/(5.39)(4.12) ???

OpenStudy (anonymous):

in degrees?

OpenStudy (amtran_bus):

My calc is in radians, and thats what she said.

OpenStudy (anonymous):

yes but I recommend you working with the exact numbers.

OpenStudy (anonymous):

but this way its right

OpenStudy (amtran_bus):

oh, I should not have rounded my answers from c^2? ok. Thanks!!!!!!!!!!!!!!!!!! So anytime you use dot product (when do you know you need to) you use in arccos and not arctan?

OpenStudy (anonymous):

yes because of the definition of the dot product, you need to remember that definition well, it has a trigonometric function embedded to it, therefore also an angle, if you search the angle, you need to use that trigonometric function that is in the formula. In case of the dot product that's the cos function.

OpenStudy (amtran_bus):

When do you know to use dot product? Thank again.

OpenStudy (anonymous):

When they ask for angles Or if two vectors are normal to each other Or maybe an engineer would use it to see from which direction and vector comes, relative to a line/plane. But now I am leaving my comfort zone (x

OpenStudy (amtran_bus):

God bless.

OpenStudy (anonymous):

in highschool/college it's 90% the first case :D